# When Unsupervised Domain Adaptation meets One-class Anomaly Detection: Addressing the Two-fold Unsupervised Curse by Leveraging Anomaly Scarcity

### v2

@@ 1 Introduction
Anomaly Detection (AD) can be seen as the identification of outliers deviating from a usual pattern. The growing interest for AD in both academia and industry is mainly due to its relevance in numerous practical scenarios, such as early disease detection in medical imaging [17, 1] and industrial inspection [39, 25, 8, 3, 24]. By definition, anomalies rarely occur. Annotating anomalous data is, therefore, often difficult and costly [16, 39, 4], hindering the collection of large-scale datasets. As a result, state-of-the-art methods mostly tackle AD as an unsupervised problem [13, 34], where the objective is to learn from the normal class.
@@
Despite achieving promising results, recent approaches in AD [33, 8, 40, 38, 16] typically assume that training and inference data are drawn from the same distribution. This assumption does not always hold in unconstrained scenarios, where a domain shift [29] between training and testing data can naturally arise due to varying setups, such as different lighting conditions and variations in object pose [4]. As a result, a model trained on a dataset sampled from a given domain, usually called source dataset, will show degraded performance when tested on a dataset from a different domain, generally termed target dataset. For instance, an AD model for medical imaging trained on images acquired using a given Medical Resonance Imaging (MRI) device can fail to generalize to samples captured with a different MRI system.
@@
To reduce such a domain gap while avoiding costly annotation efforts, Unsupervised Domain Adaptation (UDA) [42, 18] has proven to be an effective solution in binary and multi-class classification tasks [36, 18]. UDA aims at learning domain-invariant features by relying on labeled source and unlabeled target data at the same time. However, the task of unsupervised domain adaptation for unsupervised anomaly detection (UAD) is ill-posed as the goal is to: align the source and the target feature distributions using only normal source data and unlabeled target data formed by both normal and anomalous samples (see Figure 2 (c)). Hence, a direct extension of standard UDA techniques developed for binary/multi-class classification [18, 42] would not be applicable, as these methods usually aim at minimizing the distance between the estimated distributions from the entire source and target training sets. Indeed, this would lead to the erroneous alignment of both normal and anomalous target samples with normal source samples, as illustrated in Figure 1 (b). Given the learned decision boundary, this would lead to the confusion of normal and abnormal samples from the target set. As it requires addressing two unsupervised tasks simultaneously, we refer to this described problem as the two-fold unsupervised curse.
@@
To the best of our knowledge, no prior work has tried to address this two-fold unsupervised challenge, i.e., UDA for one-class image anomaly detection described in Figure 2 (c). Indeed, related works have mainly simplified the problem by either (1) assuming the availability of labeled abnormal and normal source data, resulting in UDA for a binary classification setting [19] (see Figure 2 (a)), or (2) maintaining the source one-class setup while accessing only few normal target data referred to as few-shot supervised adaptation for unsupervised anomaly detection [19, 7, 22, 23, 43] (see Figure 2 (b)). Nevertheless, annotating a few samples might still be constraining, particularly in the field of anomaly detection, where expert knowledge is often needed, such as for tumor annotation in medical images [17, 1] or for industrial inspection [8, 40, 24]. Moreover, few-shot approaches are known to be prone to overfitting issues since few shots cannot fully represent the normal target distribution [37]. This calls for a fully unsupervised domain adaptation approach that leverages the diversity of the available large unlabeled target datasets.
@@
In this paper, we propose solving the two-fold unsupervised curse by leveraging the fact that the occurrence of anomalies tends to be rare. We herein propose the first unsupervised domain adaptation framework for unsupervised image anomaly detection. Our solution starts by identifying a dominant cluster assumed to be formed by normal target data and then aligning it with normal source samples. Specifically, our method utilizes a trainable ResNet-based [15] feature extractor to process both the source and target features. A frozen CLIP visual encoder [30] is also used to generate corresponding target features, which are then clustered using K-means to identify the samples of the dominant cluster. These samples are mapped into the ResNet-based [15] feature space and aligned with the source features. For the domain adaptation task, a contrastive strategy [30, 26] ensures the similarity between the dominant target cluster and normal source samples, while for the anomaly detection task, a Deep Support Vector Data Description (DSVDD) [33] objective enforces feature compactness on the normal source data. Our framework is modular, allowing for flexible component changes, and supports various adaptation strategies, including statistical and adversarial alignment. Experiments performed on standard UDA benchmarks [35, 41, 28] for semantic anomaly detection [38] demonstrate its effectiveness. Our approach achieves state-of-the-art (SoA) performance, even compared to few-shot methods.
@@
  • A UDA method for one-class anomaly detection is introduced, leveraging a Vision Language Model, namely CLIP [30], for dominant cluster identification and alignment using a contrastive strategy.
  • Extensive experiments and analysis are conducted on several benchmarks [35, 41, 28, 20], demonstrating the relevance of the proposed framework under both fully unsupervised and few-shot settings.
@@ 2 Related Works: Anomaly detection under domain shift
Figure 2: Comparison of our setting with previous works: (a) supervised source anomaly detection with supervised domain adaptation [19], (b) unsupervised one-class source anomaly detection with few-shot domain adaptation [22, 43, 23], (c) our considered setting: unsupervised one-class source anomaly detection with unsupervised domain adaptation.
@@
Unsupervised image anomaly detection is a well-established research area [13, 34, 16, 33, 8, 40, 38] where the aim is to learn a function $\zeta$ using a single class corresponding to normal data from the normal-only dataset $\mathcal{D}^{n}=\{(\mathbf{X}_{i},y_{i});y_{i}=0\}_{i=1}^{N}$, to classify whether an input image $\mathbf{X}$ is normal ($y=0$) or not ($y=1$). This is achieved by optimizing the following objective,
$\min_{\zeta}\,\mathbb{E}_{(\mathbf{X}_{i},y_{i})\sim\mathcal{D}^{n}}\left[\mathcal{L}\left(\zeta(\mathbf{X}_{i}),y_{i}=0\right)\right],\vspace{-1mm}$ (1)
@@
where $\mathcal{L}$ is a loss enforcing feature compactness as in DSVDD [33] or a reconstruction loss typically used in autoencoders-based methods [8, 40]. Although achieving impressive performance on standard benchmarks, the majority of AD methods [13, 34, 16, 33, 8, 40, 38] overlook the domain gap problem where training and testing data denoted as $\mathcal{D}^{s}$ and $\mathcal{D}^{t}$, respectively, follow different distributions due to uncontrolled variations in the acquisition setting [4, 5]. This domain shift induces, therefore, a significant drop in performance. To solve this issue, a handful of Domain Generalization (DG) methods for UAD has been proposed recently [6, 5, 4]. Cohen et al. [6] propose a domain-disentanglement approach that removes predefined nuisance attributes (e.g., pose, lighting) from the source features using contrastive loss, preventing these factors from interfering with the anomaly task, improving the performance on unseen domains. However, without an actual target set, this method requires defining and labeling nuisance factors within the source dataset, which is challenging, as mentioned in their paper. In [5], multiple source domains are considered for learning domain-invariant features, thereby assuming the availability of diverse large-scale datasets, which is not always guaranteed. To avoid relying on multiple domains during training, a self-supervised strategy is adopted in [4]. Nevertheless, the success of this approach heavily depends on the similarity between the augmented data and target samples. As a result, it necessitates tailoring augmentation techniques to unseen target datasets, if at all possible. Given its effectiveness, domain adaptation has also been explored to address the domain shift problem in AD [7, 21, 19, 22, 23]. Those techniques usually adopt a few-shot adaptation paradigm by having access to a limited number of annotated target samples. While these methods offer innovative solutions for aligning source and target normal data, they still rely on costly annotations [16] and are exposed to overfitting risks [37]. This emphasizes the need for a fully unsupervised domain adaptation for UAD. However, solving this challenge remains challenging given the doubly unsupervised nature of the problem resulting from both UAD and UDA, which is further described in the next section.
@@ 3 The Two-fold Unsupervised Curse
Let us denote as $\mathcal{D}^{s}=\{(\mathbf{X}_{i}^{s},y_{i}^{s})\}_{i=1}^{N_{s}}$ a labeled dataset from a given domain called source formed by $N_{s}$ samples, where a sample $\mathbf{X}_{i}^{s}\in\mathbb{R}^{h\times w\times c}$ and its associated label $y_{i}^{s}\in\{0,1\}$, $\forall i=\{1,...,N_{s}\}$. Let $\mathcal{D}^{t}$ be a second unlabeled dataset from a different domain, i.e., target, denoted as $\mathcal{D}^{t}=\{\mathbf{X}_{i}^{t}\}_{i=1}^{N_{t}}$ and formed by $N_{t}$ samples where $\mathbf{X}_{i}^{t}\in\mathbb{R}^{h\times w\times c}$, $\forall i=\{1,...,N_{t}\}$. In the following, we assume that $\mathcal{D}^{t}$ shares the same label space as $\mathcal{D}^{s}$ and that there exists a domain gap between $\mathcal{D}^{s}$ and $\mathcal{D}^{t}$. The goal of Unsupervised Domain Adaptation (UDA) for anomaly detection (whether formulated as a binary or one-class classification problem), is to learn a model $\zeta:\mathbb{R}^{h\times w\times c}\rightarrow\{0,1\}$ using both $\mathcal{D}^{s}$ and $\mathcal{D}^{t}$ that generalizes to the target domain. In other words, it aims at learning a domain invariant feature extractor $f:\mathbb{R}^{h\times w\times c}\mapsto\mathcal{X}$ such that $\zeta=g\circ f$ with $g:\mathcal{X}\mapsto\{0,1\}$ being the classifier and $\mathcal{X}$ the feature space given by $f$. This objective is achieved by minimizing the following adaptation upper bound [2],
@@
While strategies for minimizing this upper bound are feasible in the context of binary or even multi-class classification [18, 36, 42], the non-availability of anomalous data during training makes it difficult in the context of one-class classification, where $d(f(\mathcal{D}^{s}),f(\mathcal{D}^{t}))$ cannot be estimated. In fact, we can only use a subset $\mathcal{D}^{s,n}\subset\mathcal{D}^{s}$ formed by normal data for training. For that reason, existing works on domain adaptation for one-class anomaly detection [7, 22, 43] revisit the formulation given in Eq (2) by slightly simplifying the problem. They pose it as a few-shot domain adaptation setting (instead of a fully unsupervised scenario). This means that they assume having access to a small subset $\mathcal{D}^{t,n}\subset\mathcal{D}^{t}$ composed of normal samples only. As a result, they reformulate Eq (2) as follows,
@@ 4 Rare Anomalies to the Rescue
For an unlabeled target dataset $\mathcal{D}^{t}=\mathcal{D}^{t,n}\cup\mathcal{D}^{t,a}$, we assume that the number of anomalous samples is significantly smaller than the number of normal samples, i.e., $|\mathcal{D}^{a}|<<|\mathcal{D}^{n}|$, with |.||.|| . | refers to the cardinality.
@@ 5 Proposed Solution to the Two-fold Unsupervised Curse
Our approach has two branches. The upper branch depicts a trainable backbone $f$ that learns from both source and target domain data. The source features are optimized using a Deep Support Vector Data Description (DSVDD) objective [33]. The lower branch focuses on visual feature extraction from the unlabeled target domain, through a frozen CLIP visual encoder [30], defined as the $\psi$ feature extractor. Clustering is applied to these visual features to estimate the dominant cluster $\mathcal{C}$. Samples identified within $\mathcal{C}$ in the $\psi$ visual encoder’s representation space are then selected within the space of the feature extractor $f$ and then aligned with the normal source features. An algorithm is provided in Section A of the Supplementary.
@@
where $\boldsymbol{\mu}^{s,n}$ is the mean of the source features. For clustering, we use a $K$-means algorithm. Note that $\psi$ can be $f$ itself or any frozen visual encoder such as CLIP [30] or DINO-v2 [27]. The dominant cluster is identified as,
@@ 6 Experimental Results
$\pm$±std $\pm$8.55 $\pm$7.55 $\pm$5.79 $\pm$9.08 $\pm$ 9.80
Table 2: Ten-run average and standard deviation of AUC (%) on the on the VisDA dataset [28] against previous SoA.

### v3

@@ 1 Introduction
Anomaly Detection (AD) can be seen as the identification of outliers deviating from a usual pattern. The growing interest for AD in both academia and industry is mainly due to its relevance in numerous practical scenarios, such as early disease detection in medical imaging [medical-application-2024-2, medical-application2024] and industrial inspection [ctt-da-fink-2024, uad-ts-2024, rd4ad-2022, warehouse-application-2018, indus-survey-2024]. By definition, anomalies rarely occur. Annotating anomalous data is, therefore, often difficult and costly [romain-2025, ctt-da-fink-2024, ad-shift-2023], hindering the collection of large-scale datasets. As a result, state-of-the-art methods mostly tackle AD as an unsupervised problem [ad-bench-2022, ruff-survey-2021], where the objective is to learn from the normal class.
@@
Despite achieving promising results, recent approaches in AD [dsvdd-2018, rd4ad-2022, rd4ad++-2023, semantic-ad-2024, romain-2025] typically assume that training and inference data are drawn from the same distribution. This assumption does not always hold in unconstrained scenarios, where a domain shift [domain-shift] between training and testing data can naturally arise due to varying setups, such as different lighting conditions and variations in object pose [ad-shift-2023]. As a result, a model trained on a dataset sampled from a given domain, usually called source dataset, will show degraded performance when tested on a dataset from a different domain, generally termed target dataset. For instance, an AD model for medical imaging trained on images acquired using a given Medical Resonance Imaging (MRI) device can fail to generalize to samples captured with a different MRI system.
@@
To reduce such a domain gap while avoiding costly annotation efforts, Unsupervised Domain Adaptation (UDA) [uda-survey-2020, uda-bench-2024] has proven to be an effective solution in binary and multi-class classification tasks [inder-wacv-2024, uda-bench-2024]. UDA aims at learning domain-invariant features by relying on labeled source and unlabeled target data at the same time. However, the task of unsupervised domain adaptation for unsupervised anomaly detection (UAD) is ill-posed as the goal is to: align the source and the target feature distributions using only normal source data and unlabeled target data formed by both normal and anomalous samples (see Figure 2 (c)). Hence, a direct extension of standard UDA techniques developed for binary/multi-class classification [uda-bench-2024, uda-survey-2020] would not be applicable, as these methods usually aim at minimizing the distance between the estimated distributions from the entire source and target training sets. Indeed, this would lead to the erroneous alignment of both normal and anomalous target samples with normal source samples, as illustrated in Figure 1 (b). Given the learned decision boundary, this would lead to the confusion of normal and abnormal samples from the target set. As it requires addressing two unsupervised tasks simultaneously, we refer to this described problem as the two-fold unsupervised curse.
@@
To the best of our knowledge, no prior work has tried to address this two-fold unsupervised challenge, i.e., UDA for one-class image anomaly detection described in Figure 2 (c). Indeed, related works have mainly simplified the problem by either (1) assuming the availability of labeled abnormal and normal source data, resulting in UDA for a binary classification setting [ildr-2019] (see Figure 2 (a)), or (2) maintaining the source one-class setup while accessing only few normal target data referred to as few-shot supervised adaptation for unsupervised anomaly detection [ildr-2019, biost-2019, msra-2021, iris-attack-detection-2022, irad-2023] (see Figure 2 (b)). Nevertheless, annotating a few samples might still be constraining, particularly in the field of anomaly detection, where expert knowledge is often needed, such as for tumor annotation in medical images [medical-application-2024-2, medical-application2024] or for industrial inspection [rd4ad-2022, rd4ad++-2023, indus-survey-2024]. Moreover, few-shot approaches are known to be prone to overfitting issues since few shots cannot fully represent the normal target distribution [challenges-few-shot-learning-2023]. This calls for a fully unsupervised domain adaptation approach that leverages the diversity of the available large unlabeled target datasets.
@@
In this paper, we propose solving the two-fold unsupervised curse by leveraging the fact that the occurrence of anomalies tends to be rare. We herein propose the first unsupervised domain adaptation framework for unsupervised image anomaly detection. Our solution starts by identifying a dominant cluster assumed to be formed by normal target data and then aligning it with normal source samples. Specifically, our method utilizes a trainable ResNet-based [resnet-2016] feature extractor to process both the source and target features. A frozen CLIP visual encoder [clip-2021] is also used to generate corresponding target features, which are then clustered using K-means to identify the samples of the dominant cluster. These samples are mapped into the ResNet-based [resnet-2016] feature space and aligned with the source features. For the domain adaptation task, a contrastive strategy [clip-2021, infonce-loss-2018] ensures the similarity between the dominant target cluster and normal source samples, while for the anomaly detection task, a Deep Support Vector Data Description (DSVDD) [dsvdd-2018] objective enforces feature compactness on the normal source data. Our framework is modular, allowing for flexible component changes, and supports various adaptation strategies, including statistical and adversarial alignment. Experiments performed on standard UDA benchmarks [office31, officehome, visda-2017] for semantic anomaly detection [semantic-ad-2024] demonstrate its effectiveness. Our approach achieves state-of-the-art (SoA) performance, even compared to few-shot methods.
@@
  • A UDA method for one-class anomaly detection is introduced, leveraging a Vision Language Model, namely CLIP [clip-2021], for dominant cluster identification and alignment using a contrastive strategy.
  • Extensive experiments and analysis are conducted on several benchmarks [office31, officehome, visda-2017, pacs-2017], demonstrating the relevance of the proposed framework under both fully unsupervised and few-shot settings.
@@ 2 Related Works: Anomaly detection under domain shift
Figure 2: Comparison of our setting with previous works: (a) supervised source anomaly detection with supervised domain adaptation [ildr-2019], (b) unsupervised one-class source anomaly detection with few-shot domain adaptation [msra-2021, irad-2023, iris-attack-detection-2022], (c) our considered setting: unsupervised one-class source anomaly detection with unsupervised domain adaptation.
@@
Unsupervised image anomaly detection is a well-established research area [ad-bench-2022, ruff-survey-2021, romain-2025, dsvdd-2018, rd4ad-2022, rd4ad++-2023, semantic-ad-2024] where the aim is to learn a function $\zeta$ using a single class corresponding to normal data from the normal-only dataset $\mathcal{D}^{n}=\{(\mathbf{X}_{i},y_{i});y_{i}=0\}_{i=1}^{N}$, to classify whether an input image $\mathbf{X}$ is normal ($y=0$) or not ($y=1$). This is achieved by optimizing the following objective,
$\min_{\zeta}\,\mathbb{E}_{(\mathbf{X}_{i},y_{i})\sim\mathcal{D}^{n}}\left[\mathcal{L}\left(\zeta(\mathbf{X}_{i}),y_{i}=0\right)\right],\vskip-2.84526pt$ (1)
@@
where $\mathcal{L}$ is a loss enforcing feature compactness as in DSVDD [dsvdd-2018] or a reconstruction loss typically used in autoencoders-based methods [rd4ad-2022, rd4ad++-2023]. Although achieving impressive performance on standard benchmarks, the majority of AD methods [ad-bench-2022, ruff-survey-2021, romain-2025, dsvdd-2018, rd4ad-2022, rd4ad++-2023, semantic-ad-2024] overlook the domain gap problem where training and testing data denoted as $\mathcal{D}^{s}$ and $\mathcal{D}^{t}$, respectively, follow different distributions due to uncontrolled variations in the acquisition setting [ad-shift-2023, ad-shift-neurips-2024]. This domain shift induces, therefore, a significant drop in performance. To solve this issue, a handful of Domain Generalization (DG) methods for UAD has been proposed recently [red-pandas, ad-shift-neurips-2024, ad-shift-2023]. Cohen et al. [red-pandas] propose a domain-disentanglement approach that removes predefined nuisance attributes (e.g., pose, lighting) from the source features using contrastive loss, preventing these factors from interfering with the anomaly task, improving the performance on unseen domains. However, without an actual target set, this method requires defining and labeling nuisance factors within the source dataset, which is challenging, as mentioned in their paper. In [ad-shift-neurips-2024], multiple source domains are considered for learning domain-invariant features, thereby assuming the availability of diverse large-scale datasets, which is not always guaranteed. To avoid relying on multiple domains during training, a self-supervised strategy is adopted in [ad-shift-2023]. Nevertheless, the success of this approach heavily depends on the similarity between the augmented data and target samples. As a result, it necessitates tailoring augmentation techniques to unseen target datasets, if at all possible. Given its effectiveness, domain adaptation has also been explored to address the domain shift problem in AD [biost-2019, tsa-2021, ildr-2019, msra-2021, iris-attack-detection-2022]. Those techniques usually adopt a few-shot adaptation paradigm by having access to a limited number of annotated target samples. While these methods offer innovative solutions for aligning source and target normal data, they still rely on costly annotations [romain-2025] and are exposed to overfitting risks [challenges-few-shot-learning-2023]. This emphasizes the need for a fully unsupervised domain adaptation for UAD. However, solving this challenge remains challenging given the doubly unsupervised nature of the problem resulting from both UAD and UDA, which is further described in the next section.
@@ 3 The Two-fold Unsupervised Curse
Let us denote as $\mathcal{D}^{s}=\{(\mathbf{X}_{i}^{s},y_{i}^{s})\}_{i=1}^{N_{s}}$ a labeled dataset from a given domain called source formed by $N_{s}$ samples, where a sample $\mathbf{X}_{i}^{s}\in\mathbb{R}^{h\times w\times c}$ and its associated label $y_{i}^{s}\in\{0,1\}$, $\forall i=\{1,...,N_{s}\}$. Let $\mathcal{D}^{t}$ be a second unlabeled dataset from a different domain, i.e., target, denoted as $\mathcal{D}^{t}=\{\mathbf{X}_{i}^{t}\}_{i=1}^{N_{t}}$ and formed by $N_{t}$ samples where $\mathbf{X}_{i}^{t}\in\mathbb{R}^{h\times w\times c}$, $\forall i=\{1,...,N_{t}\}$. In the following, we assume that $\mathcal{D}^{t}$ shares the same label space as $\mathcal{D}^{s}$ and that there exists a domain gap between $\mathcal{D}^{s}$ and $\mathcal{D}^{t}$. The goal of Unsupervised Domain Adaptation (UDA) for anomaly detection (whether formulated as a binary or one-class classification problem), is to learn a model $\zeta:\mathbb{R}^{h\times w\times c}\rightarrow\{0,1\}$ using both $\mathcal{D}^{s}$ and $\mathcal{D}^{t}$ that generalizes to the target domain. In other words, it aims at learning a domain invariant feature extractor $f:\mathbb{R}^{h\times w\times c}\mapsto\mathcal{X}$ such that $\zeta=g\circ f$ with $g:\mathcal{X}\mapsto\{0,1\}$ being the classifier and $\mathcal{X}$ the feature space given by $f$. This objective is achieved by minimizing the following adaptation upper bound [da-gen-bound-2006],
@@
While strategies for minimizing this upper bound are feasible in the context of binary or even multi-class classification [uda-bench-2024, inder-wacv-2024, uda-survey-2020], the non-availability of anomalous data during training makes it difficult in the context of one-class classification, where $d(f(\mathcal{D}^{s}),f(\mathcal{D}^{t}))$ cannot be estimated. In fact, we can only use a subset $\mathcal{D}^{s,n}\subset\mathcal{D}^{s}$ formed by normal data for training. For that reason, existing works on domain adaptation for one-class anomaly detection [biost-2019, msra-2021, irad-2023] revisit the formulation given in Eq (2) by slightly simplifying the problem. They pose it as a few-shot domain adaptation setting (instead of a fully unsupervised scenario). This means that they assume having access to a small subset $\mathcal{D}^{t,n}\subset\mathcal{D}^{t}$ composed of normal samples only. As a result, they reformulate Eq (2) as follows,
@@ 4 Rare Anomalies to the Rescue
For an unlabeled target dataset $\mathcal{D}^{t}=\mathcal{D}^{t,n}\cup\mathcal{D}^{t,a}$, we assume that the number of anomalous samples is significantly smaller than the number of normal samples, i.e., $|\mathcal{D}^{a}|<<|\mathcal{D}^{n}|$, with $|.|$ refers to the cardinality.
@@ 5 Proposed Solution to the Two-fold Unsupervised Curse
Our approach has two branches. The upper branch depicts a trainable backbone $f$ that learns from both source and target domain data. The source features are optimized using a Deep Support Vector Data Description (DSVDD) objective [dsvdd-2018]. The lower branch focuses on visual feature extraction from the unlabeled target domain, through a frozen CLIP visual encoder [clip-2021], defined as the $\psi$ feature extractor. Clustering is applied to these visual features to estimate the dominant cluster $\mathcal{C}$. Samples identified within $\mathcal{C}$ in the $\psi$ visual encoder’s representation space are then selected within the space of the feature extractor $f$ and then aligned with the normal source features. An algorithm is provided in Section A of the Supplementary.
@@
where $\boldsymbol{\mu}^{s,n}$ is the mean of the source features. For clustering, we use a $K$-means algorithm. Note that $\psi$ can be $f$ itself or any frozen visual encoder such as CLIP [clip-2021] or DINO-v2 [dinov2-2023]. The dominant cluster is identified as,
@@ 6 Experimental Results
[…]
Table 2: Ten-run average and standard deviation of AUC (%) on the on the VisDA dataset [visda-2017] against previous SoA.